1/2+(2/3 / 3/4)-(4/5 * 5/6)
Answer by Maths68(1474)

1/2+(2/3 / 3/4)-(4/5 * 5/6)
=1/2+(2/3 / 3/4)-(4/5 * 5/6)
=1/2+(2/3*4/3)-(20/30)
=1/2+(8/9)-(20/30)
=1/2+(8/9)-(2/3)
Taking LCM
=[(1*9)+(8*2)-(2*6)]/18
=(9+16-12)/18
=(9+4)/18
=13/18